# Pipeline Gradient-based Model Training on Analog In-memory Accelerators

### v1

@@ 4.2 Convergence rate of pipeline training algorithms
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}[\|\nabla f(W_{k})\|^{2}]\leq 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max}^{2}/\tau^{2}}+\sigma^{2}S$
where $S$ denotes the amplification factor given by $S:=\frac{W_{\max}/\tau^{2}}{1-W_{\max}/\tau^{2}}$.
The proof of Theorem 1 is deferred to Appendix D. Similar to the convergence of digital SGD [30], Analog SGD converges in the rate $O(\sqrt{{\sigma^{2}}/({BK})})$. However, because of the asymmetric bias, the convergence rate slows down by a factor $\frac{1}{1-W_{\max}^{2}/\tau^{2}}$, which reduces to $1$ when $\tau$ tends to infinity. Besides, the asymmetric bias introduces an additional term $\sigma^{2}S$, which is independent of $K$ and hence referred to as asymptotic error. When the AIMC accelerators can be well-designed, $\tau$ can be sufficiently large, and it does not dominate the error when $K$. Empirically, researchers also proposed some Analog SGD variants like Tiki-Taka [9, 11] to eliminate the negative impact of asymmetric bias. Next, we provide the iteration complexity of Analog SGD with asynchronous pipeline training.
Theorem 2 (Convergence rate, asynchronous pipeline).
[…]
The proof of Theorem 2 is deferred to Appendix E. Because of the stale signals, the convergence rate of the asynchronous pipeline slows down. The coefficient of the dominant term degrades from $\frac{1}{1-W_{\max}/\tau^{2}}$ to $\frac{1}{1-(1+u)W_{\max}/\tau^{2}}$, and an extra error term $O\left(\frac{1+1/u}{K}\right)$ is introduced. Fortunately, $O\left(\frac{1+1/u}{K}\right)$ is not the dominant term and we can specify $u$ as a small constant so the impact is minor.
@@
Figure 4: Training ResNet10 on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Right) Accuracy-vs-Clock cycle. (Bottom) Speedup and test accuracy. The speedup is computed by comparing the clock cycle to reach 93% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
Remark 2.
For digital SGD with the synchronous pipeline training, the convergence rate is given in [30]. We also provide the convergence analysis for SGD with asynchronous pipeline training on digital devices, detailed in Theorem 2 in the Appendix F. Unlike analog training, both synchronous and asynchronous pipeline training do not slow down digital SGD.
4.3 Sampling complexity and clock cycle complexity
The convergence rate alone does not indicate the total training time, as different algorithms require varying numbers of gradient updates. The bottleneck during a micro-batch process is the forward and backward computations, which depend on the number of micro-batches. Thus, we focus on the sample complexity of the algorithms, defined as the minimal gradient computation needed to achieve the desired error. Due to the existence of asymptotic error, we define the sample complexity of synchronous and asynchronous pipelines are the minimum micro-batch needed for achieving $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(W_{k})\|^{2}\right]\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e. $\sigma^{2}S$ for vanilla model parallelism without pipeline/synchronous pipeline and $\sigma^{2}S^{\prime}$ for asynchronous pipeline, respectively. Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, and omitting the constant $\frac{1}{1-W_{\max}^{2}/\tau^{2}}$ or $\frac{1}{1-(1+u)W_{\max}^{2}/\tau^{2}}$, we have the following corollary.
Corollary 1 (Sample complexity).
The sample complexities of model parallelism without pipeline and synchronous pipeline are $O\left(\sigma^{2}\varepsilon^{-2}\right)$, while that of the asynchronous pipeline is $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$.
Corollary 1 claims that three model parallelisms share the same dominant term $O(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while asynchronous pipeline slows down by a factor and a non-dominant term $O(\varepsilon^{-1})$. Dividing the sample complexity by the computation density, we know the number of clock cycles they need to achieve the $\varepsilon$ accuracy.
Corollary 2 (Clock cycle complexity).
The clock cycle complexities of vanilla model parallelism without pipeline, synchronous pipeline, and asynchronous pipeline are $O\left(M\sigma^{2}\varepsilon^{-2}\right)$, $O\left((M+B-1)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$, and $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$, respectively.
The results are summarized in Table 1. By overlapping the computation among micro-batches, synchronous pipeline achieves $M$ times acceleration compared to its non-pipeline counterpart without any cost. Asynchronous pipeline overlaps more computation, achieving higher computation density with slight scarification of convergence rate. When the mini-batch is split into sufficient micro-batch so that $B\gg M$, the bubble time can be omitted, and synchronous pipeline achieves the best clock cycle complexity. However, when the number of mico-batch is small, the asynchronous pipeline is able to achieve better acceleration.
@@ 4.3 Sampling complexity and clock cycle complexity
Figure 5: Training ResNet10 on CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Right) Accuracy-vs-Clock cycle. (Bottom) Speedup and test accuracy. The speedup is computed by comparing the clock cycle to reach 72% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
5 Numerical Simulations
In this section, we test our methods in an open-source AIHWKIT (AIHWKIT[31]; see github.com/IBM/aihwkit. In the simulations, we train a ResNet10 to perform an image classification task on CIFAR10 and CIFAR100 datasets; see Figure 4 and 5. The network consists of $M=6$ stages, consisting 1 convolution layer, 4 residual blocks, and 1 fully-connected layer. We map the first 5 stages on digital devices and the last one onto an AIMC (analog) device. We use digital SGD, implemented by PyTorch  library. The pipeline parallelism is also simulated on a single GPU. To get the equivalent clock cycle in the figures, we map the epoch to the clock cycle by dividing the computation density, as discussed in Section 3. Each simulation is repeated three times, and the mean and standard deviation are reported. More details are reported in Appendix G.
[…]
@@ 5 Numerical Simulations
Table 2: The training of Resnet10 model on CIFAR10 dataset. The test accuracy and standard deviation under different micro-batch sizes at clock cycle 300 (mid) and epoch 300 (end). The speedup is computed by comparing the clock cycle needed to reach 93% target accuracy. The accuracy of synchronous is not affected by micro-batch size.
Effect of micro-batch size. We also report the effect of micro-batch sizes in Table 2. We fit the mini-batch size as $B_{\text{mini}}=128$ and change the micro-batch size. For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog devices. Moreover, asynchronous pipeline training consistently results in higher speedup compared to the synchronous version. The trade-off is that the asynchronous pipeline sacrifices the accuracy and can only reach lower accuracy when the training time is sufficiently long.
@@
Effect of stage/device number. We extensively explored the impact of increasing device count on the speedup achieved through pipeline parallelism; see Figure 1. In the simulations, we train Resnet34 model on CIFAR10 dataset, which could be separated into $M=1,2,4,8,12,16$ stages. Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed. The results reveal that the asynchronous pipeline achieves linear-speedup at least within the range of 1-8 devices. Even though some latency, like communication latency, are ignored in the simulations, the comparison is still fair since the asynchronous pipeline does not introduce obvious extra latency.
6 Conclusions and Limitations
@@
This paper study the pipeline algorithms for AIMC accelerators in the convergence and computation aspects. We study the existing synchronous pipeline algorithm and propose the first asynchronous pipeline algorithm on AIMC accelerators. The synchronous pipeline achieves the best convergence rate in terms of sample complexity, but training bubbles reduce computation density. To enable more devices to be active, an asynchronous pipeline discards the synchronization at the end of one mini-batch, achieving a higher computation density, with the cost of slightly increased sample complexity. Simulations reveal the extraordinary speedup potential of our asynchronous pipeline for analog computing. A limitation of this work is that AIMC accelerator behaviors can only be simulated but not on the actual analog chips. But this is a necessary step for every new technology, and future work should verify its efficiency as the technology matures.
Supplementary Material for “Pipeline Gradient-based Model Training on Analog In-memory Accelerator”
Appendix A Analog pipeline SGD from worker and data perspectives
@@
Involving multiple devices and data at each time step, pipeline algorithms, especially asynchronous pipeline, are usually hard to figure out. To facilitate the understanding, we list the Analog SGD under two different perspectives. In worker perspective, we pay more attention to how worker $m$ processes a flow of data; see Algorithm 1. In data perspective, we focus on how a sample $\xi_{k}$ flows through each worker; see Algorithm 2.
Algorithm 1 Analog SGD with asynchronous pipeline (Worker Perspective)
@@
2:// Periphery device (worker 0) at the beginning
3:     Sample $\xi_{k}=(x_{k},y_{k})$ and Send $x^{(0)}_{k}=x_{k}$ to worker $1$
4:// Worker $m=1,2,\cdots,M$ in parallel
5:     Receive input ${\tilde{x}}^{(m)}_{k}$ from worker $m-1$
@@
9:     Receive backward error $\tilde{\delta}^{(m+1)}_{k-(M-m)}$
@@
14:     Send $\tilde{\delta}^{(M)}_{k}$ back to worker $M$
@@
Appendix B Bounds of delay errors
In the asynchronous pipeline, the computations of both forward and backward passes involve stale weights, which induces delay error on the signals. This section provides a series of the bounds of the delay error.

### v2

@@ III-C Proof sketch of Theorem 2
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]-f(W_{k})\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}$ (21)
$\ +\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}-\frac{1}{4\alpha}{\mathcal{O}}\left(\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]\right)+C_{\psi}\mathbb{E}_{\xi_{k}}[\psi_{k}]$
where $\lesssim$ performs non-essential approximation on the upper bound of descent.
From Lemma 4, we notice that the accumulated delays $\psi_{k}$ appear in the descent lemma, but intuitively they vanish when the algorithm converges according to its definition.
(Step IV) Construction of the Lyapunov function. To compensate for this term in equation (21), we design the following Lyapunov function.
${\mathbb{V}}_{k}:=f(W_{k})-f^{*}+C_{\psi}\Psi_{k}$ (22)
where the auxiliary function $\Psi_{k}$ is defined as
$\Psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (23)
and $C_{\psi}$ is a constant. To achieve the dynamics of $\Psi_{k}$, we derive that
$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (24)
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
Let the index of the inner summation of the first term in the RHS of (24) begin from $k-(M-m^{\prime})$ instead of $k+1-(M-m^{\prime})$, we have
$\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (25)
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}+M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
$=$ $\ \Psi_{k}+M\|W_{k+1}-W_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
where the last equality holds by definition of $\Delta^{(m^{\prime})}_{k}$, i.e.,
$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=$ $\ \sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}.$ (26)
Plugging (25) into (24) yields $\Psi_{k+1}\leq\Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$, which gives a negative $\psi_{k}$. We then show that the Lyapunov functions have sufficient descent if $C_{\psi}$ is properly chosen, i.e.,
$\mathbb{E}_{\xi_{k}}[{\mathbb{V}}_{k+1}]-{\mathbb{V}}_{k}\lesssim$ $\ -\frac{\alpha}{2}\left(1-(1+u)\frac{\|W_{k}\|^{2}_{\infty}}{\tau^{2}}\right)\|\nabla f(W_{k})\|^{2}+\alpha^{2}L_{f}\sigma^{2}+\frac{\alpha\sigma^{2}}{2\tau^{2}}(1+u)\|W_{k}\|^{2}_{\infty}.$ (27)
Reorganizing the inequality above yields Theorem 2.
Figure 4: Training ResNet models on CIFAR10 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). (Left) Accuracy-vs-Epoch reflects sample complexity. (Middle) Accuracy-vs-Clock cycle. (Right) Speedup of the proposed pipeline training methods on AIMC accelerators as the number of stages increases compared with the single machine training case (stage=1). We ignore the communication latency since the asynchronous pipeline does not introduce obvious extra latency.
III-D Sampling and wall-clock complexity
Sample complexity measures the computational cost required to achieve the desired accuracy, whereas in model parallelism, the computation of different stages overlaps. Therefore, we are also interested in the wall-clock time it takes. In this section, we first study the sample complexity, defined as the minimum number of gradient computations required to achieve the desired error $\epsilon$. Due to asymptotic error, we define the sample complexity in analog training as the minimum micro-batch needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e. $\sigma^{2}S$ for Analog-SGD-WOP/Analog-SGD-SP and $\sigma^{2}S^{\prime}$ for Analog-SGD-AP, respectively. Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, we have the following conclusion.
Corollary 1 (Sample complexity).
The sample complexities of Analog-SGD-WOP or Analog-SGD-SP are ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}\right)$, while that of the Analog-SGD-AP is ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$.
This shows that three model parallelisms share the same dominant term ${\mathcal{O}}(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while Analog-SGD-AP shows down by a non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$. Since the computing is overlapped, the computation density is also needed to yield the number of clock cycles needed to achieve the $\varepsilon$ accuracy.
Computation density. In an ideal scenario, a micro-batch requires $2M$ clock cycles for forward and backward passes, distributed equally across $M$ accelerators, taking 2 clock cycles per micro-batch. Computation density measures how closely an algorithm approaches this ideal. Let $N_{\text{c}}(t)$ be the number of micro-batches processed in $t$ clock cycles. We define computation density as the limit ratio of ideal processing time to real time: $\lim_{t\to\infty}\frac{2N_{\text{c}}(t)}{t}$, where the factor $2$ indicates that micro-batch processing involves forward and backward passes. Analog-SGD-WOP takes $2M$ cycles per micro-batch, yielding computation density $1/M$. In Analog-SGD-SP, it takes 2(M+B1)2(M+B-1) clock cycles to process $B$ micro-batches, resulting in a density of B/(M+B1)B/(M+B-1). Analog-SGD-AP achieves a density of $1$, as all accelerators remain active during training. Similar to sample complexity, wall-clock complexity is defined as the minimum clock cycle needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error. Wall-clock complexity is achieved by dividing the sample complexity by the computation density. The following corollary summarizes the clock cycle complexity of different pipeline strategies.
Corollary 2 (Wall-clock complexity).
The wall-clock complexities of Analog-SGD-WOP, Analog-SGD-SP, and Analog-SGD-AP are ${\mathcal{O}}\left(M\sigma^{2}\varepsilon^{-2}\right)$, ${\mathcal{O}}\left((M+B-1)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$, and ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$, respectively.
Takeaway. Analog-SGD-AP maximizes the computation density of DNN training only at the cost of a higher-order error term compared with the synchronous pipeline.
CIFAR10
@@ III-D Sampling and wall-clock complexity
CIFAR100
@@
TABLE II: Training ResNet10 on CIFAR10/CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). The speedup is computed by comparing the number of clock cycles required to achieve 93% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
@@
TABLE III: Training Resnet10 model on CIFAR10 dataset under different micro-batch size setting. Test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported. The speedup is computed by comparing the number of clock cycles required to achieve 93% target accuracy. The accuracy of synchronous is not affected by micro-batch size.
IV Numerical Simulations
In this section, we aim to numerically verify the theoretical benefits of Analog-SGD-AP. We test the analog methods in an open-source AIHWKIT (AIHWKIT[38]; see github.com/IBM/aihwkit. The pipeline parallelism is also simulated on a single GPU. To get the equivalent clock cycle in the figures, we map the epoch to clock cycle by dividing the computation density.
Simulation settings. In the simulations, we train ResNet models on two benchmark datasets: CIFAR10 and CIFAR100. The CIFAR10 dataset consists of 60,000 $\times32$ color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds. The CIFAR100 dataset is similar in format but includes 100 classes. Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default. The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage numbers $M=12,16$, the initial learning rate is $\alpha=0.01$. The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training. The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs. To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [39]. The other setting follows the default settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set to 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in [40]. A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
Sample complexity and wall-clock complexity. We train a ResNet10 for an image classification task on CIFAR10. We map the first 5 stages on digital accelerators while the last one onto an AIMC (analog) accelerator. The results on the left of Fig. 4 show that the accuracy-epoch curves of all three pipeline strategies on AIMC accelerators almost overlap and even match the performance of digital SGD, implying all of these algorithms have the same sample complexity and align with Theorem 12 and Corollary 1. Moreover, as indicated by Corollary 2, the walk-clock complexity of the Analog-SGD-AP pipeline is the lowest owing to the highest computation density. The results on the CIFAR100 training are reported in Table III, whose conclusion is the same as above.
@@ IV Numerical Simulations
Effect of stage/accelerator number. We also explored the impact of increasing the number of accelerators on the speedup achieved through pipeline parallelism; see Figure 4, Right. In the simulations, we train a ResNet34 model [41] on the CIFAR10 dataset. The model is separated into $M=1,2,4,8,12,16$ stages. Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed. The results reveal that the asynchronous pipeline achieves linear speedup at least within the range of 181-8 accelerators.
Effect of micro-batch size. We also report the effect of micro-batch sizes in Table III. We fit the mini-batch size as $B_{\text{mini}}=128$ and change the micro-batch size. For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog accelerators. Moreover, asynchronous pipeline training consistently results in higher speedup compared to the synchronous version. The trade-off is that the asynchronous pipeline sacrifices accuracy and can only reach lower accuracy when the training time is sufficiently long.
V Conclusions and Limitations
@@
This paper theoretically investigates the pipeline algorithms for AIMC accelerators with respect to convergence and computation complexity. To unlock the full potential of pipeline training, asynchronous pipeline discards the synchronization at the end of one mini-batch to maximize the accelerator usage. We theoretically demonstrate that Analog-SGD-AP achieves maximal computational density with only a slight sacrifice in sample complexity through higher-order terms, which brings extraordinary speedup on wall-clock complexity. Simulations validate our theoretical findings. A limitation of this work is the lack of real hardware validation, but this is a necessary step for every new technology and is widely used in current research regarding the algorithm development for AIMC accelerators. Future work should verify its efficiency as the technology becomes mature.
Appendix
Appendix A Analog SGD with Asynchronous Pipeline in Different Perspective
@@
Involving multiple accelerators and data at each time step, pipeline algorithms, especially the asynchronous pipeline, are usually hard to figure out. To facilitate the understanding, we list the Analog SGD under two different perspectives. In accelerator perspective, we pay more attention to how accelerator $m$ processes a flow of data; see Algorithm 1. In data perspective, we focus on how a sample $\xi_{k}$ flows through each accelerator; see Algorithm 2. To make the relationship clearer, the computation dependency is plotted in Fig. 3.
Algorithm 1 Analog SGD with asynchronous pipeline (accelerator Perspective)
@@
2:// Periphery device at the beginning
3:  Sample $\xi_{k}=(x_{k},y_{k})$ and Send $x^{(0)}_{k}=x_{k}$ to accelerator 11
4:// Accelerator $m=1,2,\cdots,M$ in parallel
5:  Receive input ${\tilde{x}}^{(m)}_{k}$ from accelerator m1m-1
@@
9:  Receive backward signal $\tilde{\delta}^{(m+1)}_{k-(M-m)}$
@@
14:  Send $\tilde{\delta}^{(M)}_{k}$ back to accelerator $M$
@@
Appendix B Useful Lemmas and their proof
Lemma 5 (Delay error of signals, restatement of Lemma 3).
Under Assumption 24, the difference between the latest and stale signals generated by Analog-SGD-AP are bounded, i.e., $\forall m\in[M+1]$
$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$, (28)
$\|{\tilde{z}}^{(m)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\sum_{m^{\prime}=1}^{m}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$, (29)
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$, (30)
where the constants above are defined by
$C_{x}^{(m)}:=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1},$ (31)
$C_{z}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{m-1},$ (32)
Proof of Lemma 3 and Lemma 5.
We first bound the delay of the forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$. For other layers $m>1$
$=$ $\ L_{g,0}\|W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}-W^{(m)}_{k}x^{(m)}_{k}\|$
where (a) follows the triangle inequality and the consistency of $\ell_{2}$-matrix norm, i.e. $\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrices and vectors, respectively. Using triangle inequality over the first term of RHS in (34), we have
and consequently,
$\|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m}\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$. (36)
$\ \|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$ (37)
where the last inequality reindexes $m^{\prime}$ in the first summation.
Noticing that (34)-(37) also implies that
$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|$ (40)
$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
$\ +W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$=$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|+(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
$\ \|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ (41)
$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
where (a) comes from Assumption 3 and Assumption 4.
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (42)
$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
Substituting (41) and (42) back into (40) we have
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (43)
where the constant $C_{\delta}^{(m)}$ is defined by
This completes the proof. ∎
Lemma 6 (Delayed error of gradient).
Denote the stale gradient as
$\tilde{\nabla}_{W^{(m)}}^{k}f:=\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$. (45)
The gradient delay is bounded by
$\|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|\leq$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|$. (46)
where the constant $C_{f}^{(m)}$ is defined by
Consequently, the norm square of the gradient delay is bounded by
$\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}\leq$ $\ C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (48)
where the constant $C_{f}$ is defined by
$C_{f}:=M\sqrt{\sum_{m=1}^{M}(C_{f}^{(m)})^{2}}.$ (49)
Proof of Lemma 6.
By definition, the delay error is bounded by
$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|\|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\delta^{(m)}_{k}\otimes x^{(m)}_{k}\|$ (50)
$=$ $\ \|(\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k})\otimes{\tilde{x}}^{(m)}_{k}+\delta^{(m)}_{k}\otimes({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|\leq\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}\|+\|\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$.
According to Lemma 8 and Lemma 5, it follows that
$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (51)
$\leq$ $\ ((L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)})\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$.
It completes the proof. ∎
Bounds of delay errors. In Analog-SGD-AP, the computations in both the forward and backward passes involve stale weights, which introduce delay error in the signals. This section presents bounds on the delay error.